OpenStudy (the_rebel_outlaw):

Find the solutions of the system y = –2x – 1 and y = 3–x–1. What are the x-coordinates of the solutions? Choose all that apply. 1 0 -1 -2

OpenStudy (jiteshmeghwal9):

y=-2x-1....(1) y=3-x-1 y=2-x ......(2)

OpenStudy (the_rebel_outlaw):

i choose the two negative numbers

OpenStudy (jiteshmeghwal9):

from eqn (2) x=2-y ...(3) substitute eqn (3) in eqn (1) to get y. &finally x=2-y

OpenStudy (the_rebel_outlaw):

so then would it be 1 and -2